# An algorithm for estimating the crossing number of dense graphs, and continuous analogs of the crossing and rectilinear crossing numbers

### v2

@@ 2.2 Cut distance and graphons
We assume some familiarity with the theory of graphons and convergent sequences of dense graphs, but still include a few of the fundamental definitions for the sake of completeness. We refer the reader to the book by Lovász [29], which we follow rather closely during for the remainder of the section, for on an in-depth treatment of the subject.
Let $G=(V,E)$ be an edge weighted graph. For any two subsets $S,T\subset V$, let $E_{G}(S,T)$ denote the set of edges with one endpoint in $S$ and the other one in $T$, and let $e_{G}(S,T)$ be the total weight of the elements of $E_{G}(S,T)$, where edges with both endpoints in $S\cap T$ are counted twice.
Given a positive integer $m$, the m blow-up $G[m]$ of $G$ is the edge-weighted graph obtained by replacing each vertex $v$ of $G$ by an independent set $U_{v}$ with $m$ elements, and then setting the weight of every edge between $U_{u}$ and $U_{v}$ (with $u\neq v$) to be $w_{G}(u,v)$. For any partition $\mathcal{P}=\{V_{1},V_{2},\dots,V_{n}\}$ of $V$, let $G_{\mathcal{P}}$ denote the edge weighted graph with vertex set $V$ such that
@@
The labeled cut distance between two edge weighted graphs $G_{1}$ and $G_{2}$ on the same finite vertex set $V$ is defined as
@@
If $G_{1}$ and $G_{2}$ are are defined on possibly different $n$-element vertex sets, then
@@
where $G_{1}^{\prime}$ and $G_{2}^{\prime}$ range over all graphs with vertex set $\{1,2,\dots,n\}$ which are isomorphic to $G_{1}$ and $G_{2}$, respectively. We are ready to define the cut distance between two arbitrary graphs. Given two edge weighted graphs $G_{1}$ and $G_{2}$ on $m$ and $n$ vertices, respectively, the cut distance between is given by
@@
which can be shown to be well defined. The distance function $\delta_{\square}$ is a pseudometric on the set of edge weighted graphs (it is symmetric and satisfies the triangle inequality, but $\delta_{\square}(G_{1},G_{2})$ does not imply that $G_{1}$ and $G_{2}$ are isomorphic). We say that a sequence edge weighted graphs $G_{1},G_{2},\dots$ is convergent if it is Cauchy with respect to $\delta_{\square}$.
Throughout this paper, measurability is always with respect to the Lebesgue $\sigma$-algebra, although working with Borel sets would not make a significant difference. We denote the Lebesgue measure by $\lambda$. A kernel is a symmetric measurable function $W:[0,1]^{2}\rightarrow\mathbb{R}$ (by symmetric, we mean that $W(x,y)=W(y,x)$), the space of all kernels is denoted by $\mathcal{W}$. A graphon is a kernel whose image is a subset of $[0,1]$, we denote the space of graphons by $\mathcal{W}_{0}$.
@@
For every graph $G$ with vertices $v_{1},v_{2},\dots,v_{n}$, the graphon $W_{G}$ is constructed by splitting $[0,1]$ into $n$ intervals $I_{1},I_{2},\dots,I_{n}$ of measure $1/n$ and setting $W_{G}(x,y)=w_{G}(v_{i},v_{j})$ for all $x\in I_{i},y\in I_{j}$. To be precise, $W_{G}$ depends on both $G$ and an ordering of its vertices, but this will not be an issue, since any two such graphons that arise from the same graph are weakly isomorphic. Given any two edge weighted graphs $G_{1}$ and $G_{2}$, we have that $\delta_{\square}(G_{1},G_{2})=\delta_{\square}(W_{G_{1}},W_{G_{2}})$. For every convergent graph sequence $G_{1},G_{2},\dots$, the graphons $W_{G_{1}},W_{G_{2}},\dots$ converge, with respect to $\delta_{\square}$, to some graphon $W$ (or, rather, to the class of $W$ in $\widetilde{\mathcal{W}_{0}}$), and we say that the graph sequence converges to $W$. Conversely, for every graphon $W$ there exists a convergent graph sequence $G_{1},G_{2},\dots$ such that $W_{G_{1}},W_{G_{2}},\dots$ converges to $W$ with respect to $d_{\square}$.
@@
whenever $x\in S_{i}$ and $y\in S_{j}$. Let $\mathcal{P}_{1},\mathcal{P}_{2},\dots$ be a sequence of partitions of $[0,1]$ into measurable sets such that each pair of points of $[0,1]$ is separated by all but a finite number of partitions of the sequence. Then, as $n$ goes to infinity, $W_{\mathcal{P}_{n}}$ converges to $W$ almost everywhere for every $W\in\mathcal{W}$.
@@ 2.3 Estimable parameters
By a graph parameter, we mean a function that assigns a real number to each graphs (or edge weighted graph) and is constant on each isomorphism class. A graph parameter $f$ is said to be estimable if there is another graph parameter $g$, which we call a test parameter for $f$, with the following property: For every $\varepsilon>0$ there exist an integer $k$ such that if $G=(V,E)$ is a graph on at least $k$ vertices and $X$ is a random $k$-element subset of $V$, then the probability that
$|f[X]-g(G[X])|>\varepsilon$
is at most $\varepsilon$ (here, $G[X]$ denotes the subgraph of $G$ that is induced by $X$). It is no hard to see that if $f$ is estimable then can always use $g=f$ (cf. [21, 9]). As shown by Borgs et al. [9], the following properties are equivalent for every graph parameter $f$.
@@
If $f$ is estimable, then the functional $\hat{f}$ mentioned in (c) also satisfies $\lim_{n\rightarrow\infty}f(G_{n})=f(W)$ whenever $G_{1},G_{2},\dots$ converges to $W$ and the number of vertices of $G_{n}$ goes to infinity with $n$. We often refer to a functional on $\mathcal{W}_{0}$ as a graphon parameter. The authors of [9] also proved that the following three properties together are equivalent to $f$ being estimable.
@@
    For every $\varepsilon>0$ there exists an $\varepsilon^{\prime}$ such that any two graphs $G_{1}$ and $G_{2}$ on the same vertex set with $d_{\square}(G_{1},G_{2})<\varepsilon^{\prime}$ satisfy $|f(G_{1})-f(G_{2})|<\varepsilon$.
@@
    $|f(G)-f(G\cup K_{1})|\rightarrow 0$ as the number of vertices of $G$ goes to infinity.
@@ 2.4 The Frieze-Kannan regularity lemma
According to Szemerédi’s regularity lemma [40], for every $\varepsilon>0$ there exists an $M(\varepsilon)$ such that for every graph we can find an equitable partition of its vertices into no more than $M(\varepsilon)$ parts with the property that all but at most an $\varepsilon$ fraction of the pairs of parts are $\varepsilon$-regular. Szemerédi’s regularity lemma is one of the most powerful tools in the study of dense graphs; unfortunately, it is not very practical for algorithmic purposes, since $M(\varepsilon)$ grows extremely fast as $\varepsilon$ goes to 00 (see [22] for details about the asymptotic behaviour of the optimal value of $M(\varepsilon)$). As was done by the authors of [18], we circumvent this issue by means of a variant of the regularity lemma developed by Frieze and Kannan [19]. Given a graph $G=(V,E)$, we say that an equitable partition $\mathcal{P}=\{V_{1},V_{2},\dots,V_{n}\}$ of $V$ is a Frieze-Kannan $\varepsilon$-regular partition if $d_{\square}(G,G_{\mathcal{P}})\leq\varepsilon$. The algorithmic version of the Frieze-Kannan regularity lemma stated below is due to Dellamonica et al. [15].
@@
There exist a deterministic algorithm and an absolute constant $c$ such that that for any $\varepsilon>0$ and any $n$-vertex graph $G=(V,E)$, computes a Frieze-Kannan $\varepsilon$-regular partition of $G$ with no more than $2^{\varepsilon^{-c}}$ classes in $2^{2^{\varepsilon^{-c}}}n^{2}$-time.
@@ 2.5 Cycle separators for planar graphs
The planar separator theorem [42, 28] states that for that every $n$-vertex planar graph $G=(V,E)$ there exists a partition $V=A\sqcup B\sqcup C$ of its vertex set such that $|B|=O(\sqrt{n})$, $|A|,|C|\leq 2n/3$, and there are no edges between $A$ and $C$. This result is a cornerstone of the study of planar graphs, and several generalizations and variants of it have been discovered throughout the years. We will make use of the following version, due to Miller [32].
@@
Let $G$ be an embedded triangulation [^5] By a triangulation we mean a maximal planar graph. Every face of an embedded triangulation has precisely three vertices on its boundary. with a non-negative weight assigned to each vertex so that the total sum of the weights is $1$. Then, there exists a simple cycle of length at most $L\sqrt{n}$ in $G$ such that the vertices that lie in its interior have total weight at most $2/3$ and the same is true for those that belong to its exterior. Here, $L$ is an absolute positive constant.
@@ 3 Estimating the crossing number
We say that a drawing $\mathcal{D}$ of an edge weighted graph $G$ is locally optimal if the value of $\mathop{\mathrm{cr}}(G,\mathcal{D})$ can not be reduced by erasing a single curve that represents an edge and then redrawing the edge in some other way. If there are no edges of weight 00, then any locally optimal drawing is necessarily simple. We remark, however, that not every locally optimal drawing attains the crossing number of the graph. This section begins with a partitioning result about simple and locally optimal graph drawings, which is somewhat similar in spirit to the classical cutting lemma for line arrangements [^6] The cutting lemma is a powerful tool in both discrete and computational geometry. In its two dimensional version, it can be stated as follows: Let $\mathcal{L}$ be a finite family of lines on the plane and let $\varepsilon>0$. Then it is possible to subdivide the plane into $O(1/\varepsilon^{2})$ convex regions, none of which is crossed by more than $\varepsilon|\mathcal{L}|$ lines of $\mathcal{L}$. See [5, 10] for various proofs of this result and its higher dimensional variants, as well as several of their applications., and will play a crucial role in our proof that any two graphs which are close with respect to $d_{\square}$ have similar crossing numbers (Theorem 3.3).
@@
Let $G$ be an $n$-vertex graph and let $\mathcal{D}$ be a simple and locally optimal drawing of $G$. Then, for any $\varepsilon\in[0,1]$ the plane can be subdivided into $O(1/\varepsilon^{2})$ closed, connected and interior disjoint regions, each of which has the following properties:
@@
    Any vertex and any other point within it can be connected by a simple curve that lies completely within the region, and whose relative interior, which does not go through any vertex or crossing point of $\mathcal{D}$, has a total of no more than $\varepsilon n^{2}$ intersection points with the curves that represent the edges of $G$ that have no endpoint in that same region.
@@
If $\varepsilon\leq n^{-1/2}$, the subdivision can be obtained by splitting the plane into $n+1$ regions, $n$ of which are very small and contain precisely one vertex each. For $\varepsilon\geq 1$ we can use a single region, so from now on we assume that $n^{-1/2}<\varepsilon<1$. A simple curve whose relative interior does not pass through any vertex or crossing point of $\mathcal{D}$ will be called clean.
@@
We begin subdividing the plane into $O(1/\varepsilon^{2})$ closed, connected and interior disjoint regions which contain no more than $\lceil\varepsilon^{2}n\rceil$ vertices of $G$ each; this will be achieved through repeated applications of Theorem 2.3. More precisely, we will build a sequence $R_{0},R_{1},\dots,R_{t}$ of regions as follows: Let $R_{0}$ denote the entire plane. By adding a dummy vertex at each crossing point, $\mathcal{D}$ gives rise to an embedded planar graph, which can then be transformed into an embedded triangulation $T_{0}$ by adding edges within each face with more than three vertices on its boundary. Furthermore, this can be done so that the degree in $T_{0}$ of every vertex of $G$ is upper bounded by $3n$. Assign a weight of $1/n$ to every vertex of $G$ and a weight of 00 to every dummy vertex and apply Theorem 2.3 to $T_{0}$, thus obtaining a cycle $C_{0}$ of length less than $Ln^{2}$ whose interior and exterior each contain at most $32n/$ 3 vertices of $G$. Every vertex of $G$ that belongs to $C_{0}$ can be assigned to either the interior or the exterior of the cycle so that the number of vertices in any of these two regions plus the number of vertices that have been assigned to it is still no more than $32n/$ 3. By modifying the cycle as depicted in Figure 1, we obtain a $C_{0}^{\prime}$ with less than $Ln^{2}+n\cdot 3n$ nodes which contains no vertex of $G$. Denote the two regions induced by this curve as $R_{1}$ and $R_{2}$; these contain at most $32n/$ 3 vertices of $G$ each. For every $i$, $n_{i}$ will be used to denote the number of vertices of $G$ that lie in $R_{i}$. Now, we iteratively apply the procedure described below to subdivide each $R_{i}$ that contains $n_{i}>\lceil\varepsilon^{2}n\rceil$ vertices of $G$ and has not yet been subdivided into smaller regions.
@@
If we restrict our attention to the boundary of $R_{i}$ and the portion of $\mathcal{D}$ that lies inside this region, we get a planar graph. By adding edges, this graph can be made into an embedded triangulation $T_{i}$ so that, for every connected component of the complement of $R_{i}$, all the edges that are drawn inside it are incident to the same vertex. Furthermore, the process we are about to discuss will ensure that no vertex of $G$ lies on the boundary of any of the $R_{j}$’s so, again, we can assume that the degree of every vertex of $G$ in $T_{i}$ is upper bounded by $3n$. If we assign a weight of $1/n_{i}$ to each of the vertices of $G$ that belong to $T_{i}$, then Theorem 2.3 yields a cycle $C_{i}$ of length at most $L\sqrt{m_{i}}$, where $m_{i}$ denotes the order of $T_{i}$. From now on, we will ignore the sections of $C_{i}$ that are not contained in $R_{i}$; note that there is at most one of these for every connected component of the complement of $R_{i}$. As we did with $C_{0}$, $C_{i}$ can be modified slightly to obtain a curve (or, rather, a collection of curves) with no more than $L\sqrt{m_{i}}+3n_{i}$ nodes that does not go through any vertex of $G$ and such that both its exterior and interior contain at most $32n_{i}/$ 3 vertices of $G$. One could try to use this curve directly to split $R_{i}$ into smaller parts, but this could cause the number of regions to blow up in the case that the curve shares many points with the boundary of $R_{i}$. Thus, we first modify $C_{i}$ as shown in Figure 2 for every contiguous sequence of nodes and arcs that it shares with the boundary of $R_{i}$ and connects two arcs that belong to the interior of $R_{i}$. This results in a new $C_{i}^{\prime}$ with at most $4(L\sqrt{m_{i}}+3n_{i})$ nodes, which is then used to split $R_{i}$ into two smaller connected regions. After enough steps, we arrive at a subdivision of the plane into closed, connected and interior disjoint regions that contain at most $\lceil\varepsilon^{2}n\rceil$ vertices of $G$ each. These regions will be referred to as fundamental regions from now on.
@@
Figure 1: A small portion of $C_{0}$ has been highlighted in red. Assuming that the vertex $v$ of $G$ has been assigned to the shaded region, we alter the cycle around a small neighborhood of $v$ so that this point belongs to the interior of the the shaded region defined by $C_{0}^{\prime}$. This increases the number of nodes of the cycle by less than the degree of $v$ in $T_{0}$.

### v4

@@ 2.2 Cut distance and graphons
We expect the reader to be somewhat familiar with the theory of graphons and convergent sequences of dense graphs, but include the required fundamental definitions for the sake of completeness. We refer the reader to the book by Lovász [34], which we follow rather closely during for the remainder of the section, for an in-depth treatment of the subject.
Let $G=(V,E)$ be an edge-weighted graph. For any two subsets $S,T\subset V$, let $E_{G}(S,T)$ denote the set of edges with one endpoint in $S$ and the other one in $T$, and let $e_{G}(S,T)$ be the total weight of the elements of $E_{G}(S,T)$, where edges with both endpoints in $S\cap T$ are counted twice.
Given a positive integer $m$, the m blow-up $G[m]$ of $G$ is the edge-weighted graph obtained by replacing each vertex $v$ of $G$ by an independent set $U_{v}$ with $m$ elements, and then setting the weight of every edge between $U_{u}$ and $U_{v}$ (with $u\neq v$) to be $w_{G}(u,v)$. For any partition $\mathcal{P}=\{V_{1},V_{2},\dots,V_{n}\}$ of $V$, let $G_{\mathcal{P}}$ denote the edge-weighted graph with vertex set $V$ such that
@@
The labeled cut distance between two edge-weighted graphs $G_{1}$ and $G_{2}$ on the same finite vertex set $V$ is defined as
@@
If $G_{1}$ and $G_{2}$ are are defined on possibly different $n$-element vertex sets, then we write
@@
where $G_{1}^{\prime}$ and $G_{2}^{\prime}$ range over all graphs with vertex set $\{1,2,\dots,n\}$ which are isomorphic to $G_{1}$ and $G_{2}$, respectively. We are now ready to define the cut distance between two arbitrary graphs. Given two edge-weighted graphs $G_{1}$ and $G_{2}$ on $m$ and $n$ vertices, respectively, the cut distance between them is given by
@@
which can be shown to be well defined. The distance function $\delta_{\square}$ is a pseudometric on the set of edge-weighted graphs (it is symmetric and satisfies the triangle inequality, but $\delta_{\square}(G_{1},G_{2})=0$ does not imply that $G_{1}$ and $G_{2}$ are isomorphic). We say that a sequence edge-weighted graphs $G_{1},G_{2},\dots$ is convergent if it is Cauchy with respect to $\delta_{\square}$.
Throughout this paper, measurability is always considered with respect to the Lebesgue $\sigma$-algebra, although working with Borel measurability would not make a significant difference. We denote the Lebesgue measure by $\lambda$. A kernel is a symmetric measurable function $W:[0,1]^{2}\rightarrow\mathbb{R}$ (by symmetric, we mean that $W(x,y)=W(y,x)$); the space of all kernels is denoted by $\mathcal{W}$. A graphon is a kernel whose image is a subset of $[0,1]$, and we denote the space of graphons by $\mathcal{W}_{0}$.
@@
For every graph $G$ with vertices $v_{1},v_{2},\dots,v_{n}$, the graphon $W_{G}$ is constructed by splitting $[0,1]$ into $n$ intervals $I_{1},I_{2},\dots,I_{n}$ of measure $1/n$ and setting $W_{G}(x,y)=w_{G}(v_{i},v_{j})$ for all $x\in I_{i},y\in I_{j}$. To be precise, $W_{G}$ depends on both $G$ and an ordering of its vertices, but this will not be an issue, since any two such graphons that arise from the same graph are weakly isomorphic. Given any two edge-weighted graphs $G_{1}$ and $G_{2}$, we have that $\delta_{\square}(G_{1},G_{2})=\delta_{\square}(W_{G_{1}},W_{G_{2}})$. For every convergent graph sequence $G_{1},G_{2},\dots$, the graphons $W_{G_{1}},W_{G_{2}},\dots$ converge, with respect to $\delta_{\square}$, to some graphon $W$ (or, rather, to the class of $W$ in $\widetilde{\mathcal{W}_{0}}$), and we say that the graph sequence converges to $W$. Conversely, for every graphon $W$ there exists a convergent graph sequence $G_{1},G_{2},\dots$ such that $W_{G_{1}},W_{G_{2}},\dots$ converges to $W$ with respect to $\delta_{\square}$.
@@
whenever $x\in S_{i}$ and $y\in S_{j}$. Let $\mathcal{P}_{1},\mathcal{P}_{2},\dots$ be a sequence of partitions of $[0,1]$ into measurable sets such that each pair of points of $[0,1]$ lie in different parts for all but a finite number of partitions of the sequence. Then, as $n$ goes to infinity, $W_{\mathcal{P}_{n}}$ converges to $W$ almost everywhere for every $W\in\mathcal{W}$.
@@ 2.3 Estimable parameters
By a graph parameter, we mean a function that assigns a real number to each graph (or edge-weighted graph) and is constant on each isomorphism class. A graph parameter $f$ is said to be estimable if there is another graph parameter $g$, which we call a test parameter for $f$, with the following property: For every $\varepsilon>0$ there exist an integer $k$ such that if $G=(V,E)$ is a graph on at least $k$ vertices and $X$ is a random $k$-element subset of $V$, then
$\mathbb{P}[|f[G]-g(G[X])|>\varepsilon]\leq\varepsilon,$
where $G[X]$ denotes the subgraph of $G$ that is induced by $X$. It is no hard to see that if $f$ is estimable then can always use $g=f$ (cf. [24, 9]). As shown by Borgs et al. [9], the following properties are equivalent for every graph parameter $f$.
@@
If $f$ is estimable, then the functional $\hat{f}$ mentioned in (c) also satisfies $\lim_{n\rightarrow\infty}f(G_{n})=\hat{f}(W)$ whenever $G_{1},G_{2},\dots$ converges to $W$ and the number of vertices of $G_{n}$ goes to infinity with $n$. We often refer to a functional on $\mathcal{W}_{0}$ as a graphon parameter. It was also proven in [9] that the following three properties together are equivalent to $f$ being estimable.
@@
    For every $\varepsilon>0$, there exists an $\varepsilon^{\prime}$ such that any two graphs $G_{1}$ and $G_{2}$ on the same vertex set with $d_{\square}(G_{1},G_{2})<\varepsilon^{\prime}$ satisfy $|f(G_{1})-f(G_{2})|<\varepsilon$.
@@
    $|f(G)-f(G\cup K_{1})|\rightarrow 0$ as the number of vertices of $G$ goes to infinity. Here, $G\cup K_{1}$ denotes the graph obtained by adding an isolated node to $G$.
@@ 2.4 The Frieze-Kannan regularity lemma
According to Szemerédi’s regularity lemma [48], for every $\varepsilon>0$ there exists an $M(\varepsilon)$ such that for every graph we can find an equitable partition of its vertices into no more than $M(\varepsilon)$ parts with the property that all but at most an $\varepsilon$ fraction of the pairs of parts are $\varepsilon$-regular. Szemerédi’s regularity lemma is one of the most powerful tools in the study of dense graphs; unfortunately, it is not very practical for algorithmic purposes, since $M(\varepsilon)$ grows extremely fast as $\varepsilon$ goes to 00 (see [25] for details on the asymptotic behavior of the optimal value of $M(\varepsilon)$). As was done by the authors of [21], we circumvent this issue by means of a variant of the regularity lemma developed by Frieze and Kannan [22]. Given a graph $G=(V,E)$, we say that an equitable partition $\mathcal{P}=\{V_{1},V_{2},\dots,V_{n}\}$ of $V$ is a Frieze-Kannan $\varepsilon$-regular partition if $d_{\square}(G,G_{\mathcal{P}})\leq\varepsilon$. The algorithmic version of the Frieze-Kannan regularity lemma stated below is due to Dellamonica et al. [17].
@@
There exist a deterministic algorithm and an absolute constant $c$ that, for any $\varepsilon>0$ and any $n$-vertex graph $G$, computes a Frieze-Kannan $\varepsilon$-regular partition of $G$ with no more than $2^{\varepsilon^{-c}}$ classes in $2^{2^{\varepsilon^{-c}}}n^{2}$-time.
@@ 2.5 Cycle separators for planar graphs
The planar separator theorem from [50, 33] states that for every $n$-vertex planar graph $G=(V,E)$ there exists a partition $V=A\sqcup B\sqcup C$ such that $|B|=O(\sqrt{n})$, $|A|,|C|\leq 2n/3$, and there are no edges between $A$ and $C$. This result is a cornerstone of the study of planar graphs, and several generalizations and variants of it have been discovered over the years. We will make use of the following version, due to Miller [37].
@@
Let $G$ be an embedded triangulation [^5] A triangulation is a maximal planar graph. By an embedded planar graph we simply mean the graph together with a drawing of it where no two edges cross. Every face of an embedded triangulation has precisely three vertices on its boundary. We remark that all ways of embedding a triangulation are, in a precise sense, combinatorially equivalent, so the theorem could have also been stated without ever referring to a particular embedding. with a non-negative weight assigned to each vertex so that the total sum of the weights is $1$. Then, there exists a simple cycle of length at most $L\sqrt{n}$ in $G$ such that the vertices that lie in its interior have total weight at most $2/3$ and the same is true for those that belong to its exterior. Here, $L$ is an absolute positive constant.
2.6 Triangulating with small degrees
In order to make use of Theorem 2.3, we need to be able to transform embedded planar graphs into embedded triangulations by adding some nodes and edges. Moreover, it will be important for our purposes that the arising triangulation does not have too many new vertices, and that degrees of its nodes are not too large. The following result, which essentially appears as Lemma 2.1 in [42], allows us to do just that.
Lemma 2.4.
Let $G$ be a connected, embedded, $n$-vertex planar graph. Suppose that every vertex of $G$ has degree at most $d$ for some $d\geq 3$. Then, the embedding can be extended to a triangulation with at most $19n$ vertices and all degrees bounded from above by $3d$.
Remark. The proof of this lemma is rather simple. In[42], the result is stated for two-connected graphs embedded in a genus $g$ surface. An inspection of the proof reveals that the two-connectedness is not required for graphs embedded in the plane.
@@ 3 Estimating the crossing number
3.1 Subdividing drawings
We say that a drawing $\mathcal{D}$ of an edge-weighted graph $G$ is locally optimal if the value of $\mathop{\mathrm{cr}}(G,\mathcal{D})$ cannot be reduced by erasing a single curve that represents an edge and then redrawing the edge in some other way. If there are no edges of weight 00, then any locally optimal drawing is necessarily simple. We remark, however, that not every locally optimal drawing attains the crossing number of the graph. This section begins with a result about subdividing simple and locally optimal graph drawings, which is somewhat similar in spirit to the classical cutting lemma for line arrangements [^6] The cutting lemma is a powerful tool in both discrete and computational geometry. In its two dimensional version, it can be stated as follows: Let $\mathcal{L}$ be a finite family of lines on the plane and let $\varepsilon>0$. Then, it is possible to subdivide the plane into $O(1/\varepsilon^{2})$ convex regions, none of which is crossed by more than $\varepsilon|\mathcal{L}|$ lines of $\mathcal{L}$. See [5, 11] for various proofs of this result and its higher dimensional variants, as well as several of their applications., and will play a crucial role in our proof that any two graphs which are close with respect to $d_{\square}$ have similar crossing numbers (see Theorem 3.3).
@@ 3.1 Subdividing drawings
Let $G$ be an $n$-vertex edge-weighted graph and let $\mathcal{D}$ be a simple and locally optimal drawing of $G$. Then, for any $\varepsilon\in(0,1)$ the plane can be subdivided into $O(1/\varepsilon^{2})$ closed, connected and interior disjoint regions, each of which has the following properties:
@@
    Any vertex of $G$ and any other point which are both contained in the region can be connected by a simple curve that lies completely within the region and does not go through any vertex or crossing, and whose relative interior has no more than $\varepsilon n^{2}$ intersection points with the curves that represent the edges of $G$ with no endpoint in that same region.
@@
If $\varepsilon\leq n^{-1/2}$, the subdivision can be obtained by splitting the plane into $n+1$ regions, $n$ of which are very small and contain precisely one vertex each. From now on, we assume that $n^{-1/2}<\varepsilon<1$.
We begin by subdividing the plane into $O(1/\varepsilon^{2})$ closed, connected and interior disjoint regions which contain no more than $\lceil\varepsilon^{2}n\rceil$ vertices of $G$ each, and whose boundaries, which do not contain any vertex of $G$, have a total of $O(n^{2}/\varepsilon)$ intersections with with the edges of $\mathcal{D}$. Here, by an intersection what we really mean is a connected component of the intersection between an edge and the boundary of one of the regions; an edge and the boundary of a region can induce multiple intersections, but the number of intersections between them is always an integer. This will be achieved through repeated applications of Theorem 2.3 and Lemma 2.6. More precisely, we will build a sequence $R_{0},R_{1},\dots,R_{t}$ of regions as follows:
@@
Let $R_{0}$ denote the entire plane. By adding a dummy vertex at each crossing point, $\mathcal{D}$ gives rise to an embedded planar graph. Using Lemma 2.6, this planar graph can then be transformed into an embedded triangulation $T_{0}$. One can readily check that the original planar graph has at most $n^{4}/4$ vertices, so $T_{0}$ has at most $19n^{4}/4<5n^{4}$ vertices, each of degree at most $3n$. Assign a weight of $1/n$ to every vertex of $G$ and a weight of 00 to every other vertex and apply Theorem 2.3 to $T_{0}$, thus obtaining a cycle $C_{0}$ of length less than $\sqrt{5}Ln^{2}$ whose interior and exterior each contain at most $32n/$ 3 vertices of $G$. Every vertex of $G$ that belongs to $C_{0}$ can be assigned to either the interior or the exterior of the cycle so that the number of vertices in any of these two regions plus the number of vertices that have been assigned to it is still no more than $32n/$ 3. By modifying the triangulation and the cycle $C_{0}$ as depicted in Figure 1, we obtain a new cycle $C_{0}^{\prime}$ with less than $\sqrt{5}Ln^{4}+n\cdot 3n$ nodes which contains no vertex of $G$. Denote the two closed regions induced by this curve as $R_{1}$ and $R_{2}$; these contain at most $32n/$ 3 vertices of $G$ each.
@@
For every $i$, we write $n_{i}$ to denote the number of vertices of $G$ that lie in $R_{i}$. Now, we iteratively subdivide each $R_{i}$ that contains $n_{i}>\lceil\varepsilon^{2}n\rceil$ vertices of $G$ and has not yet been subdivided into smaller regions; the rest of this paragraph describes how to carry out this subdivision. First, the boundary of $R_{i}$ and the portion of $\mathcal{D}$ that is contained in this region are turned into an embedded planar graph by adding a dummy node at every intersection (either between two edges or between the boundary and an edge). Let $m_{i}$ denote the order of this graph. A face of this embedded planar graph will be called null if its interior is contained in the complement of $R_{i}$. Observe that there is precisely one null face for each connected component of the complement of $R_{i}$. Once again, Lemma 2.6 allows us to transform this planar graph into an embedded triangulation with no more than $19m_{i}$ nodes, each of degree at most $3n$. Now, we modify the triangulation by deleting all nodes and edges added in the interior of each null face, and then re-triangulating each such face simply by picking one of the vertices on its boundary and connecting it to all others using curves drawn within the face. This is possible, since among the vertices on the boundary of any null face there is at least one (there are at least two, actually [^7] In general, given an embedded planar graph this will true of every face whose boundary is a simple curve. This fact might look more familiar to the reader in its equivalent, “inverted” form: For any set of pairwise interior disjoint diagonals of a convex polygon, there are at least two nodes of the polygon which are not incident to any of the diagonals.) which is adjacent to only two other vertices from the said boundary, and any such vertex can be used to draw the edges from. Let $T_{i}$ be the triangulation that is obtained in this manner. As mentioned above, the process we are about to discuss will ensure that no vertex of $G$ ever lies on the boundary of any of the $R_{j}$’s (in particular, this implies that no vertex of $G$ will ever lie on the boundary of a null face), so it is still true that the degree of every vertex of $G$ in $T_{i}$ is upper bounded by $3n$. If we assign weight $1/n_{i}$ to each of the vertices of $G$ in $R_{i}$ and weight 00 to all other vertices of $T_{i}$, then Theorem 2.3 yields a cycle $C_{i}$ of length at most $L\sqrt{19m_{i}}$, where $m_{i}$ denotes the order of $T_{i}$. As we did for $C_{0}$, $T_{i}$ and $C_{i}$ can be modified slightly to obtain a cycle $C_{i}^{\prime}$ with no more than $L\sqrt{19m_{i}}+3n_{i}$ nodes that does not go through any vertex of $G$ and such that the intersections of $R_{i}$ with both the interior and the exterior of the cycle each contain at most $32n_{i}/$ 3 vertices of $G$. From now on, we shall ignore the edges of $C_{i}^{\prime}$ that are drawn on the interior of a null face, as well as all nodes of $C_{i}^{\prime}$ which are incident to two ignored edges; note that there are at most two of these edges and one such node for each null face. One could try to use $C_{i}^{\prime}$ in order to split $R_{i}$ into smaller parts, but this may cause the number of regions to blow up in the case that $C_{i}^{\prime}$ intersects the boundary of $R_{i}$ at multiple places. Thus, we first modify $C_{i}^{\prime}$ as shown in Figure 2 for every contiguous sequence of nodes and edges that it shares with the boundary of $R_{i}$ and lies between two edges of $C_{i}^{\prime}$ that are draw inside $R_{i}$. This results in a new $C_{i}^{\prime\prime}$ with at most $4(L\sqrt{19m_{i}}+3n_{i})$ nodes, which is then used to split $R_{i}$ into two smaller connected regions with at most $32n_{i}/$ 3 vertices of $G$ each.
@@
Figure 1: A small portion of $C_{0}$ has been highlighted in red. Assuming that the vertex $v$ of $G$ has been assigned to the shaded region, we alter the triangulation and the cycle around a small neighborhood of $v$ so that this point belongs to the interior of the shaded region defined by $C_{0}^{\prime}$. This increases the number of nodes of the cycle (and the whole graph) by less than the degree of $v$ in $T_{0}$.